# Quench dynamics of the quantum XXZ chain with staggered interactions: Exact results and simulations on digital quantum computers

### v2

@@ IV.1 Hadamard test for $a_{k}$ estimation
Figure 9: Hadamard test circuit for computing the real part of the expectation value $\text{Re}(\langle{\bm{0}}|{\bm{U}}|{\bm{0}}\rangle)$, where $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$ is an $N$-qubit state and $\left|{0}\right\rangle_{a}$ is the initial state of an ancilla qubit.
@@
Our first numerical experiment estimates the coefficients (amplitudes) $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$ using the Hadamard test [39]. The Hadamard test circuit, shown in Fig. 9, consists of two Hadamard gates $\bm{H}$ acting on a single ancilla qubit, which serves as the control qubit for a controlled-$\bm{U}$ gate. The unitary is defined as $\bm{U}=\bm{U}_{\Phi_{k}}^{\dagger}\bm{U}_{\Psi_{0}}$, where $\bm{U}_{\Psi_{0}}$ and $\bm{U}_{\Phi_{k}}$ generate the states $\left|{\Psi_{0}}\right\rangle$ and $\left|{\Phi_{k}}\right\rangle$, respectively, from a register of $N$ qubits initialized to $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$:
@@
The measurement is performed on the ancilla qubit in the $Z$ basis. The probabilities of measuring 0 and $1$ in the ancilla qubit are
@@ IV.2 Time evolution circuit
While the Hadamard test provides direct amplitude estimation of time-evolved states in the Bell basis, this approach yields accurate results only for very small systems on real quantum devices. As system size increases, deeper circuits and diminishing amplitude magnitudes significantly reduce accuracy on noisy hardware, making this method impractical for $N>4$ in our simulations.
@@
in terms of Pauli gates $\bm{X}$, $\bm{Y}$ and $\bm{Z}$. The quantum circuits implementing the three exponential operators with $ZZ$, XXXX, and $YY$ couplings are illustrated in Fig. 14(a), 14(b), and 14(c), respectively. In this setup, the exponential operators are related to each other through the basis transformation: $({\bm{u}^{X}})^{\dagger}\bm{Z}\bm{u}^{X}=\bm{X}$ and $({\bm{u}^{Y}})^{\dagger}\bm{Z}\bm{u}^{Y}=\bm{Y}$ with $\bm{u}^{X}=\bm{H}$ and $\bm{u}^{Y}=\bm{HS}^{\dagger}$, where $\bm{S}$ is a phase shift gate which rotates a state vector a $\pi/2$ radian around the $Z$-axis.
@@ IV.3 Randomized measurements
The measurement protocol proceeds as follows: (i) Prepare the postquench state $\rho(t)$ of $N$ qubits at time $t$ using the time-evolution circuit. (ii) Apply a random unitary transformation, $\rho\rightarrow\bm{U}\rho\bm{U}^{\dagger}$, where $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ and each $\bm{u}_{i}$ is randomly drawn from $\{\bm{H},\bm{H}\bm{S}^{\dagger},\bm{I}\}$, thereby rotating the local measurement basis to a randomly chosen Pauli basis $\{X,Y,Z\}$. (iii) Measure each qubit in the standard $Z$ basis, obtaining a bitstring $\mathbf{b}=z_{1}\cdots z_{N}$ with $z_{i}=0,1$. For each fixed unitary $\bm{U}$, the measurement is repeated $N_{M}$ times. The full procedure is then repeated for $N_{U}$ independently sampled unitaries. The resulting $N_{U}\times N_{M}$ bitstrings and corresponding sets of single-qubit unitaries are stored for classical post-processing.
We apply two post-processing protocols to estimate the second-order Rényi entanglement entropy (EE) and the Loschmidt echo. In the first protocol, the purity of a subsystem $A$, required for the Rényi EE, is estimated from statistical correlations of randomized measurements via [40, 41, 42]:
@@
where $D[\mathbf{b}_{A},\mathbf{b}^{\prime}_{A}]\equiv\left|{\{i\in A|z_{i}\neq z^{\prime}_{i}\}}\right|$ is the Hamming distance between two bitstrings $\mathbf{b}_{A}=z_{1}\cdots z_{N_{A}}$ and $\mathbf{b}^{\prime}_{A}=z^{\prime}_{1}\cdots z^{\prime}_{N_{A}}$, and $P_{\bm{U}}(\mathbf{b}_{A})$ denotes the probability of measuring bitstring $\mathbf{b}_{A}$ under unitary $\bm{U}$. Similarly, the overlap between two $N$-qubit states $\rho_{1}$ and $\rho_{2}$ can be estimated using [42]:
@@
where $P^{(1)}_{\bm{U}}$ and $P_{\bm{U}}^{(2)}$ are the probabilities based on $\rho_{1}$ and $\rho_{2}$, respectively. From the subsystem purity, we compute the Rényi EE of a half-chain, while identifying $\rho_{1}=\rho(0)=\left|{\Psi_{0}}\right\rangle\left\langle{\Psi_{0}}\right|$ and $\rho_{2}=\rho(t)$ yields the Loschmidt echo.
@@
A second post-processing protocol employs the pairs of random unitaries and measurement outcomes, $\{(\bm{U},\mathbf{b})\}$, to construct classical shadows of the quantum state  [43, 44]. If $\left|{\mathbf{b}}\right\rangle=\left|{z_{1}\cdots z_{N}}\right\rangle$ is the measurement outcome in the $Z$ basis after applying the unitary transformation $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ to an $N$-qubit state $\rho$, a classical snapshot of the state can be constructed as $\bigotimes_{i=1}^{N}\bm{u}_{i}^{\dagger}\left|{z_{i}}\right\rangle\left\langle{z_{i}}\right|\bm{u}_{i}$. The reduced state on a subsystem $A$ is then reconstructed via inverting an $N_{A}$-qubit depolarizing channel [43, 44]:
@@
This rule, together with the identity in Eq. (60), provides an efficient and scalable method for evaluating the purity using the classical shadow formalism.
To obtain the Loschmidt echo, we first estimate the $N$-qubit state for a given set of random unitaries $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ at time $t$ by averaging over $N_{M}$ measurement outcomes:
@@
In Fig. 15 and Fig. 16, we present the results for the half-chain Rényi EE and the Loschmidt echo, respectively, comparing data obtained from random Pauli measurements with the exact solutions. The same measurement data are post-processed using Hamming distance method and classical shadow technique. According to previous studies [40, 42, 43, 44], for a fixed total number $M=N_{M}\times N_{U}$, the Hamming distance method yields more accurate estimates when the number of measurement shots $N_{M}$ is large, whereas the classical shadow technique performs better when $N_{U}$ is large. Without optimizing the ratio of $N_{M}$ and $N_{U}$ for our setting, we chose $N_{U}=2^{6}$ and $N_{M}=2^{13}$.
@@
Without any error mitigation, the simulated half-chain EE shows overall good agreement with the exact results. For the Loschmidt echo, where the full system contributes, some noticeable deviations from the exact solutions appear at certain times. Nonetheless, the results from the two post-processing methods agree well with each other across all cases studied.
@@ V Summary and discussion
We have analytically and numerically investigated quantum quenches in the XXZ chain with alternating interactions in the flat-band limit. The quench protocol we considered interchanges the odd- and even-bond strengths of an initially fully dimerized chain. By formulating the problem in the Bell basis, we established selection rules identifying all Bell states with nonzero expansion coefficients in the time-dependent state for arbitrary even system sizes. For a chain of length $N$, all nonzero coefficients have equal magnitude $\left|{a_{k}}\right|=2^{-(N/2-1)}$. We further derived exact relations connecting the dynamical half-chain entanglement entropies of the $N=4$ system to those of any even system with $N>4$. From the exact results for the Loschmidt echo up to $N=16$, we identified a series of Loschmidt zeros occurring at critical times in certain finite chains. No such zeros were found in the isotropic XXX case ($\Delta=1$) or in chains whose length is not a multiple of four. Both the entanglement entropy and the Loschmidt echo exhibit persistent oscillations, reflecting the absence of relaxation in this flat-band system. For $N>4$, these observables are periodic with period $2q\pi$ when $J\Delta=p/q$ is an irreducible fraction.
@@ Acknowledgements.
We acknowledge support from the IBM Q Hub at National Taiwan University, and the National Science and Technology Council (NSTC). Y.C.L. was supported by the NSTC under Grants No. NSTC 114-2119-M-007-013 and 113-2119-M-007-013; she is grateful to H.-C. Hsu and H.-C. Chang for useful discussions and previous collaboration. F.I. was supported by the Hungarian Scientific Research Fund under Grant No. K146736 and by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary; he also acknowledges travel support from the NSTC and National Chengchi University.
Appendix A Entanglement entropy for $n=4$ XX- and XXX-chains
In this Appendix, we provide the analytical expressions for the dynamical von Neumann EE for the $n=4$ XX- and XXX-chains, i.e. with $\Delta=0$ and $\Delta=1$.
For the $n=4$ XX chain, the eigenvalues of the reduced density matrix for a half-chain are given by
@@ Appendix A Entanglement entropy for $n=4$ XX- and XXX-chains
Here, $\lambda_{1}$ and $\lambda_{2}$ are non degenerate, $\lambda_{3}$ and $\lambda_{4}$ are both 4-fold degenerate and $\lambda_{5}$ is 6-fold degenerate. From Eq. (21), we obtain the half-chain von Neumann EE at time $t$:
@@
where $\lambda_{1}$ is 9-fold degenerate, $\lambda_{2}$ is 6-fold degenerate and $\lambda_{3}$ is non-degenerate. The half-chain von Neumann EE at time $t$ is then given by

### v3

@@ IV.1 Hadamard test for $a_{k}$ estimation
Figure 10: Hadamard test circuit for computing the real part of the expectation value $\text{Re}(\langle{\bm{0}}|{\bm{U}}|{\bm{0}}\rangle)$, where $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$ is an NN-qubit state and $\left|{0}\right\rangle_{a}$ is the initial state of an ancilla qubit.
@@
Our first numerical experiment estimates the coefficients (amplitudes) $a_{k}=\langle{\Phi_{k}}|{\Psi_{0}}\rangle$ using the Hadamard test [39]. The Hadamard test circuit, shown in Fig. 10, consists of two Hadamard gates $\bm{H}$ acting on a single ancilla qubit, which serves as the control qubit for a controlled-$\bm{U}$ gate. The unitary is defined as $\bm{U}=\bm{U}_{\Phi_{k}}^{\dagger}\bm{U}_{\Psi_{0}}$, where $\bm{U}_{\Psi_{0}}$ and $\bm{U}_{\Phi_{k}}$ generate the states $\left|{\Psi_{0}}\right\rangle$ and $\left|{\Phi_{k}}\right\rangle$, respectively, from a register of NN qubits initialized to $\left|{\bm{0}}\right\rangle=\left|{0}\right\rangle^{\otimes N}$:
@@
To implement the controlled-$\bm{U}$ gate, we first construct the circuits preparing $\left|{\Psi_{0}}\right\rangle$ and $\left|{\Phi_{k}}\right\rangle$ and convert them into unitary operators using the to_gate() function in Qiskit [69]. We then form $\bm{U}=\bm{U}_{\Phi_{k}}^{\dagger}\bm{U}_{\Psi_{0}}$ and convert it to a controlled operation,
$\bm{C}_{\bm{U}}=\begin{pmatrix}\bm{I}&0\\ 0&\bm{U}\end{pmatrix}\,$, (51)
using ControlledGate, with an ancilla qubit as the control qubit initialized by a Hadamard gate.
The ancilla qubit is measured in the $Z$ basis after a second Hadamard gate. The probabilities of measuring 0 and 11 in the ancilla qubit are
@@ IV.2 Time evolution circuit
While the Hadamard test enables direct amplitude estimation of time-evolved states in the Bell basis, it yields accurate results only for very small systems on real quantum devices. As system size grows, circuit depth increases substantially, degrading accuracy on noisy hardware. Error mitigation techniques, such as readout error mitigation, zero-noise extrapolation, and Pauli twirling, can partially compensate for hardware noise [70]; however, the exponential decay of amplitude magnitudes with system size ultimately limits the practical scalability of this approach in our simulations.
@@
in terms of Pauli gates $\bm{X}$, $\bm{Y}$ and $\bm{Z}$. The quantum circuits implementing the three exponential operators with ZZZZ, XXXX, and YYYY couplings are illustrated in Fig. 15(a), 15(b), and 15(c), respectively. In this setup, the exponential operators are related to each other through the basis transformation: $({\bm{u}^{X}})^{\dagger}\bm{Z}\bm{u}^{X}=\bm{X}$ and $({\bm{u}^{Y}})^{\dagger}\bm{Z}\bm{u}^{Y}=\bm{Y}$ with $\bm{u}^{X}=\bm{H}$ and $\bm{u}^{Y}=\bm{HS}^{\dagger}$, where $\bm{S}$ is a phase shift gate which rotates a state vector a $\pi/2$ radian around the $Z$-axis.
@@ IV.3 Randomized measurements
The measurement protocol proceeds as follows: (i) Prepare the postquench state $\rho(t)$ of NN qubits at time $t$ using the time-evolution circuit. (ii) Apply a random unitary transformation, $\rho\rightarrow\bm{U}\rho\bm{U}^{\dagger}$, where $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ and each $\bm{u}_{i}$ is randomly drawn from $\{\bm{H},\bm{H}\bm{S}^{\dagger},\bm{I}\}$, thereby rotating the local measurement basis to a randomly chosen Pauli basis $\{X,Y,Z\}$. (iii) Measure each qubit in the standard $Z$ basis, obtaining a bitstring $\mathbf{b}=z_{1}\cdots z_{N}$ with $z_{i}=0,1$. For each fixed unitary $\bm{U}$, the measurement is repeated $N_{M}$ times. The full procedure is then repeated for $N_{U}$ independently sampled unitaries. The resulting $N_{U}\times N_{M}$ bitstrings and corresponding sets of single-qubit unitaries are stored for classical post-processing.
We apply two post-processing protocols to estimate the second-order Rényi entanglement entropy ($E$) and the Loschmidt echo. In the first protocol, the purity of a subsystem AA, required for the Rényi $E$, is estimated from statistical correlations of randomized measurements via [40, 41, 42]:
@@
where $D[\mathbf{b}_{A},\mathbf{b}^{\prime}_{A}]\equiv\left|{\{i\in A|z_{i}\neq z^{\prime}_{i}\}}\right|$ is the Hamming distance between two bitstrings $\mathbf{b}_{A}=z_{1}\cdots z_{N_{A}}$ and $\mathbf{b}^{\prime}_{A}=z^{\prime}_{1}\cdots z^{\prime}_{N_{A}}$, and $P_{\bm{U}}(\mathbf{b}_{A})$ denotes the probability of measuring bitstring $\mathbf{b}_{A}$ under unitary $\bm{U}$. Similarly, the overlap between two NN-qubit states $\rho_{1}$ and $\rho_{2}$ can be estimated using [42]:
@@
where $P^{(1)}_{\bm{U}}$ and $P_{\bm{U}}^{(2)}$ are the probabilities based on $\rho_{1}$ and $\rho_{2}$, respectively. From the subsystem purity, we compute the Rényi $E$ of a half-chain, while identifying $\rho_{1}=\rho(0)=\left|{\Psi_{0}}\right\rangle\left\langle{\Psi_{0}}\right|$ and $\rho_{2}=\rho(t)$ yields the Loschmidt echo.
@@
A second post-processing protocol employs the pairs of random unitaries and measurement outcomes, $\{(\bm{U},\mathbf{b})\}$, to construct classical shadows of the quantum state  [43, 44]. If $\left|{\mathbf{b}}\right\rangle=\left|{z_{1}\cdots z_{N}}\right\rangle$ is the measurement outcome in the $Z$ basis after applying the unitary transformation $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ to an NN-qubit state $\rho$, a classical snapshot of the state can be constructed as $\bigotimes_{i=1}^{N}\bm{u}_{i}^{\dagger}\left|{z_{i}}\right\rangle\left\langle{z_{i}}\right|\bm{u}_{i}$. The reduced state on a subsystem AA is then reconstructed via inverting an $N_{A}$-qubit depolarizing channel [43, 44]:
@@
This rule, together with the identity in Eq. (61), provides an efficient and scalable method for evaluating the purity using the classical shadow formalism. This method has been incorporated into a software package for sampling and post-processing randomized measurements [72].
To obtain the Loschmidt echo, we first estimate the NN-qubit state for a given set of random unitaries $\bm{U}=\bigotimes_{i=1}^{N}\bm{u}_{i}$ at time $t$ by averaging over $N_{M}$ measurement outcomes:
@@
In Fig. 16 and Fig. 17, we present the results for the half-chain Rényi $E$ and the Loschmidt echo, respectively, comparing data obtained from random Pauli measurements with the exact solutions. The same measurement data are post-processed using Hamming distance method and classical shadow technique. According to previous studies [40, 42, 43, 44], for a fixed total number $M=N_{M}\times N_{U}$, the Hamming distance method yields more accurate estimates when the number of measurement shots $N_{M}$ is large, whereas the classical shadow technique performs better when $N_{U}$ is large. Without optimizing the ratio of $N_{M}$ and $N_{U}$ for our setting, we chose $N_{U}=2^{6}$ and $N_{M}=2^{13}$.
@@
Without any error mitigation, the simulated half-chain $E$ shows overall good agreement with the exact results. For the Loschmidt echo, where the full system contributes, some noticeable deviations from the exact solutions appear at certain times. Nonetheless, the results from the two post-processing methods agree well with each other across all cases studied.
@@ V Summary and discussion
We have analytically and numerically investigated quantum quenches in the XXZ chain with alternating interactions in the flat-band limit. The quench protocol we considered interchanges the odd- and even-bond strengths of an initially fully dimerized chain. By formulating the problem in the Bell basis, we established selection rules identifying all Bell states with nonzero expansion coefficients in the time-dependent state for arbitrary even system sizes. For a chain of length NN, all nonzero coefficients have equal magnitude $\left|{a_{k}}\right|=2^{-(N/2-1)}$. We further derived exact relations connecting the dynamical half-chain entanglement entropies of the $N=4$ system to those of any even system with $N>4$. We also derived a series of closed-form solutions for the Loschmidt echo up to $N=12$, which can be extended to numerically exact results for arbitrary NN using the exact coefficients $\left|{a_{k}}\right|$.
[…]
@@ Acknowledgements.
We are grateful to Balázs Dóra for valuable discussions. We acknowledge support from the IBM Q Hub at National Taiwan University, the National Science and Technology Council (NSTC), and the National Center for Theoretical Sciences. Y.C.L. was supported by the NSTC under Grants No. NSTC 114-2119-M-007-013 and 113-2119-M-007-013; she is grateful to H.-C. Hsu and H.-C. Chang for previous collaboration. F.I. was supported by the Hungarian Scientific Research Fund under Grant No. K146736 and by the National Research, Development and Innovation Office of Hungary (NKFIH) within the Quantum Information National Laboratory of Hungary; he also acknowledges travel support from the NSTC and National Chengchi University.
Appendix A Entanglement entropy for $n=4$ $X$- and XXX-chains
In this Appendix, we provide the analytical expressions for the dynamical von Neumann $E$ for the $n=4$ $X$- and XXX-chains, i.e. with $\Delta=0$ and $\Delta=1$.
For the $n=4$ $X$ chain, the eigenvalues of the reduced density matrix for a half-chain are given by
@@ Appendix A Entanglement entropy for $n=4$ $X$- and XXX-chains
Here, $\lambda_{1}$ and $\lambda_{2}$ are non degenerate, $\lambda_{3}$ and $\lambda_{4}$ are both 4-fold degenerate and $\lambda_{5}$ is 6-fold degenerate. From Eq. (21), we obtain the half-chain von Neumann $E$ at time $t$:
@@
where $\lambda_{1}$ is 9-fold degenerate, $\lambda_{2}$ is 6-fold degenerate and $\lambda_{3}$ is non-degenerate. The half-chain von Neumann $E$ at time $t$ is then given by
@@ Appendix B Loschmidt echo for $n=6$
Figure 18: A comparison of the Loschmidt echo for $n=6$ and various vaIlues of $\Delta$, obtained from Eqs. (69)-(71) (solid) and exact diagonalization (dotted), confirming the correctness of the analytical results. The coupling strength is fixed at $J=1$
@@
Figure 18 shows the Loschmidt echo for $n=6$ with various values of $\Delta$ obtained from Eqs. (69)-(71) compared with data from exact diagonalization, confirming the correctness of the analytical results.
Appendix C Loschmidt echo for the $X$-chain with $n=5,7$ and 99
The analytical expressions of the Loschmidt echo for the $X$-chain with $n=5,7$ and 99 are given below:
${\cal L}_{XX}^{n=5}(t)=\frac{1}{4^{8}}\left[23126+31440\cos(Jt)+9720\cos(2Jt)\right$.
$\left.+1200\cos(3Jt)+50\cos(4Jt)\right]\;,$ (72)
${\cal L}_{XX}^{n=7}(t)=\frac{1}{4^{12}}\left[5015004+7604688\cos(Jt)\right$.
$\left.+3281278\cos(2Jt)+776776\cos(3Jt)\right$.
$\left.+94276\cos(4Jt)+5096\cos(5Jt)+98\cos(6Jt)\right]\;,$ (73)
${\cal L}_{XX}^{n=9}(t)=\frac{1}{4^{16}}\left[1134397990+1826957088\cos(Jt)\right$.
$\left.+949082256\cos(2Jt)+312914784\cos(3Jt)\right$.
$\left.+63548856\cos(4Jt)+7564320\cos(5Jt)\right$.
$\left.+487152\cos(6Jt)+14688\cos(7Jt)+162\cos(8Jt)\right]\;.$ (74)